If the marginal propensity to consume (for the economy) is equal to 0.80, it must be the case that the marginal propensity to save equals:

a.

20

b.

0.20

c.

1

d.

0.90

Unlike the Classicals, Keynes held that a market-based capitalist economy was:

a.

similar to the feudal economy.

b.

inherently unstable and that small disturbances in output, prices, and employment were likely to be magnified by the invisible hand of the market.

c.

not capable of being regulated.

d.

capable of resolving its own difficulties.
Answer #1


Question 1

MPC = 0.80

As we know that,

MPS + MPC = 1

MPS = 1 - MPC

MPS = 1 - 0.80 = 0.20

Thus,

The marginal propensity to save is 0.20.

Hence, the correct answer is the option (b).

Question 2

Keynes stated that economy does not operate always at full employment. He also stated that economy can not return to its full employment on its own. It needs intervention.

This was based on his observation that capitalist economy was inherently unstable and that small disturbances in output, prices, and employment were likely to be magnified by the invisible hand of the market.

Hence, the correct answer is the option (b).